OpenStudy (anonymous):

MEDAL TO BEST ANSWER W/ EXPLANATION RATIOS AND PROPORTIONS Write and solve the proportion: 3 is to 12 as w is to 8. w equals two w equals four w equals twenty-four w equals thirty-two

OpenStudy (anonymous):

@sourwing @Reaper534

OpenStudy (anonymous):

3 is to 12 as w is to 8. w = 2

OpenStudy (anonymous):

Ok, but how did you figure that out?

OpenStudy (anonymous):

3/12 = w/8 Cross multiply: 12w = 24 w = 2

OpenStudy (anonymous):

Okay, that makes sense